In order to figure out the difference betweent simulated result and measurement, I tried to measuren the electronic noise by following ways as show in attachment 1
1.measure from the satellite box by SR785 at ETMY ,calibrate to counts by divide by 3267.8. while at that conditin, the set up is in suspension.
2. measure after ADC by diagnostics test tools, with set up on table in history and on uspension currently.
3. use the caculated butterfly channel.
the results are shown in attachmemt 2. The overall nosie level are still much higher than simulation.
If we have some data with one of the optics clamped and the open light hitting the PD, or with the OSEMs removed and sitting on the table, that would be useful for evaluating the end-to-end noise of the OSEM circuit. It seems like we probably have that due to the vent work, so please post the times here if you have them.
The ETMX OSEMs have been attached to its Satellite box and plugged in for the last 10 days or so, with the PD exposed to the unobstructed LED. I pulled the spectrum of one of the sensors (mean detrended, I assume this takes care of removing the DC value?). The DQed channels claim to record um (the raw ADC counts are multiplied by a conversion factor of 0.36). For comparison, re-converted the y-axis for the measured curve to counts, and multiplied the total noise curve from the LISO simulation by a factor of 3267.8cts/V (2^16cts/20V) so the Y axis is noise in units of counts/rtHz. At 1Hz, there is more than an order of magnitude difference between the simulation and the measurement which makes me suspect my y-axis conversion, but I think I've done this correctly. Can such a large discrepancy be solely due to thick film resistors?